what is the y=mx+b form of these systems
x+y=4
2x+y=6

Answer by drk(1908)

To write in the form y = mx + b, we need to solve for y.
(i) x+y=4
subtract x to get
y = -x+4
---
(ii) 2x+y=6
subtract 2x to get
y = -2x + 6
---
Both are now in y = mx+ b form
--
If you wanted the values of x and y that make them consistent, then we get
x = 2
y = 2